OpenStudy (anonymous):

An investor opened an account with an online brokerage firm and deposited $1,000.00 into the account. She then used all the money in the account to buy some shares in a company, costing her $15.00 in fees, and made a 10% profit on the remainder of her money. She then sold the shares, costing her $15.00 in fees, and bought some shares in another company with the remainder of her money, costing her $15.00 in fees, and she made a 10% profit on the remainder of her money. Finally, she sold the shares, costing her $15.00 in fees. Which of these recursive formulas will produce the sequence tha

OpenStudy (anonymous):

A) t1 = 1000, tn = (1.1)(tn – 1 – 15) – 15 B) t1 = 1000, tn = (1.1)(tn – 1 – 15) c) t1 = 1000, tn = (1.1)(tn – 1) – 15 D) t1 = 1000, tn = (1.1)(tn – 1) – 30

OpenStudy (anonymous):

@strawberrysorbet @Kaiden37 @iGreen

OpenStudy (anonymous):

I'm terribly sorry, I haven't studied this :(

OpenStudy (anonymous):

oh ok